OpenStudy (anonymous):

Find the vector and parametric equations for this line: goes through (7,5) with inclination 45º. need hellllppppp:/

OpenStudy (jack1):

m = 1/1 = 1 as inclination is 45 degrees do you understand that?

OpenStudy (jack1):

as you only have a single point the most advanced parametric equation you can work out is a linear equation, which takes the form y = mx +c

OpenStudy (jack1):

@ewa you following ok so far?

OpenStudy (anonymous):

perfectly

OpenStudy (jack1):

good, so you have the equation type (y = mx + c) and you have m, now tell me what c is?

OpenStudy (anonymous):

-2

OpenStudy (jack1):

spot on, so there's ur equation

OpenStudy (anonymous):

thank you

OpenStudy (jack1):

always welcome ;)

OpenStudy (phi):

this gives an explanation of the different equations http://tutorial.math.lamar.edu/Classes/CalcIII/EqnsOfLines.aspx the vector form of the line through (7,5) with inclination 45º. First, find a vector that points in the same direction as the line you want to define. V= (1,1) (with a tail at (0,0) and head at (1,1) makes an angle of 45 deg with the x-axis) your equation will be: P= (7,5) + t*(1,1) where P is any point on the line (a position vector) and "t" is a scalar (that is, a number)